Circular No. 7167

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 DN_{3} The apparently asteroidal object 1999 DN_{3}, observed by K. Korlević

and M. Jurić at Višnjan (0.41-m f/4.3 reflector + CCD) on Feb. 19.0 and 24.0 UT (MPC 33833, MPS 4018), was linked by G. V. Williams, Minor Planet Center, to observations on Apr. 6 and 14 in routine asteroidal astrometry from LINEAR. Owing to the unusual nature of the orbit, computed yesterday, the object was added to The NEO Confirmation Page. In response to this, further observations were reported on May 14.2 by D. A. Klinglesmith, III, and R. Huber (Etscorn Observatory) and by G. Hug (Farpoint Observatory). Williams also identified LONEOS observations of the object on Apr. 10. In addition, C. W. Hergenrother, Lunar and Planetary Laboratory, reports that observations made last night with the 1.5-m Catalina reflector show the object to be cometary, with a compact, well-condensed 10′′ coma and a strongly curved 30′′ tail, starting in p.a. 45^{o} and curving to p.a. 335^{o}. The discovery observation is given below; all the available astrometric observations appear on MPEC 1999-J39.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Feb. 18.97279  13^{h}02^{m}10.^{s}41  −6^{o}47′52′′.8  18.7

The following orbital elements by Williams are from 20 observations, Feb. 18–May 14 (mean residual 0′′.85):

Epoch = 1998 Sept. 24.0 TT

T = 1998 Sept.27.0993 TT  ω = 160.^{o}8131
e = 0.135067  Ω = 5.9635 }2000.0
q = 3.909738 AU  i = 18.7527

a = 4.520279 AU n^{o} = 0.1025548 P = 9.611 years

1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
May 12 12^{h}16.^{m}32 − 6^{o}39.^{′}8 3.193 3.975 135.^{o}4 10.^{o}3 18.5
22 12 14.05 − 6 47.7 3.310 3.981 125.3 12.0 18.6
June 1 12 13.41 − 7 02.9 3.442 3.987 115.7 13.3 18.7
11 12 14.34 − 7 25.7 3.584 3.993 106.5 14.1 18.8
21 12 16.75 − 7 56.0 3.734 3.999  97.7 14.6 18.9
July 1 12 20.52 − 8 33.4 3.886 4.006  89.3 14.7 19.0
11 12 25.49 − 9 17.3 4.039 4.013  81.3 14.5 19.1
21 12 31.54 −10 07.1 4.188 4.020  73.5 14.0 19.2
31 12 38.52 −11 02.2 4.332 4.027  66.0 13.3 19.2
1999 May 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green